Q4.From the given data, calculate i) GDP at

market price ii) GDP at factor cost and GNP

Household Consumption ExpenditureRs 550 billion

Government ConsumptionExpenditureRs 250 billion

Gross fixedcapital formationRs 100 billion

DepreciationRs 150 billion

Indirect TaxesRs 160 billion

SubsidiesRs 40 billion

ExportsRs 200 billion

ImportsRs 250 billion

Net incomefrom abroadRs 159 billion


Ans

GDP at market price= 850 billion

GDP at factor cost= 730 billion

GNP at factor cost=880 billion

GNP at market price=1000 billion


1
Expert's answer
2021-07-05T09:01:33-0400

Solution:

i.). GDP at market price = Household consumption expenditure + Government consumption expenditure + Gross fixed capital formation + Net exports

GDP at market price = 550 + 250 + 100 + (200 – 250) = 850 billion

 

ii.). GDP at factor cost = GDP at market price – Depreciation + Net income from abroad – Net indirect taxes

GDP at factor cost = 850 – 150 + 150 – (160 – 40) = 730 billion

 

iii.). GNP at market price = GDP at market price + Net income from abroad

GNP at market price = 850 + 150 = 1000 billion

 

iv.). GNP at factor price = GNP at market price – Net indirect taxes

GNP at factor price = 1000 – (160 – 40) = 1000 – 120 = 880 billion